# NRR-Phi: Text-to-State Mapping for Ambiguity Preservation in LLM Inference

### v2

@@ Abstract
On a test set of 68 ambiguous sentences, the resulting states preserve interpretive multiplicity: using hybrid extraction we obtain mean state entropy $H=1.087$ bits across ambiguity categories, compared to $H=0$ for collapse-based baselines that commit to a single interpretation. We additionally instantiate the rule-based conflict detector for Japanese markers (kedo, kamoshirenai, etc.) to illustrate cross-lingual portability of the conflict detection stage. This framework extends Non-Resolution Reasoning (NRR) by providing the missing algorithmic bridge between text and the NRR state space, enabling architectural collapse deferment in LLM inference.
@@ 1.2 Non-Resolution Reasoning
Non-Resolution Reasoning (NRR) [1] offers an alternative paradigm that treats ambiguity preservation as a valid computational mode. NRR introduced the state space $\mathcal{S}=\{(v_{i},c_{i},w_{i})\}$ where multiple interpretations coexist with independent activation weights, and demonstrated that “NRR-lite” implementations (k=2k{=}2 separate embeddings with context gate) maintain high entropy ($H=0.91$ bits, near-maximum $1.0$) at ambiguous turns while baseline models collapse ($H=0.15$ bits).
@@ 1.4 Paper Organization
Section 2 reviews related work on uncertainty in LLM inference. Section 3 presents the formal framework for text-to-state mapping. Section 4 provides algorithms and implementation details. Section 5 reports experimental results. Section 6 briefly illustrates application to psychological support contexts. Section 7 discusses limitations and future work. Section 8 concludes.
@@ 2.2 Multi-Hypothesis Approaches
Beam search [7] and diverse decoding [8] maintain multiple candidates during generation but ultimately select a single output. Mixture-of-Experts architectures [9, 10] route inputs to specialized subnetworks but still produce collapsed outputs.
@@ 2.4 Relation to NRR Framework
This paper builds directly on NRR-Core [1]. That work established the state space formalism and demonstrated entropy preservation through prompt engineering. NRR defines eight operators ($\sigma,\alpha,\rho,\iota,\delta,\tau,\kappa,\pi$) for manipulating non-collapsing states (see Appendix D for summary).
@@ 5.5 Results: LLM-Based Extraction
To address the limitations of rule-based extraction, we evaluated LLM-based interpretation extraction using three models: ChatGPT, Gemini, and Claude. [^1] All experiments used the free tier of each service’s web interface (January 2026). Specific version numbers are not displayed in free-tier interfaces; based on publicly documented model availability, the models were approximately: ChatGPT (GPT-5.2), Claude (Claude 3.7 Sonnet or later), and Gemini (Gemini 2.5 Flash). Each model was prompted to enumerate all possible interpretations for 28 English sentences (8 epistemic + 10 lexical + 10 structural).
@@ 5.6 Combined Results
Overall 68 2.34 1.087 Hybrid
Table 4: Combined results using hybrid extraction. All categories achieve $H>0$, with overall mean entropy $H=1.087$ bits.
@@ Appendix C Sample Outputs
Appendix D NRR Operator Summary
NRR uses eight operators to transform the state space $\mathcal{S}$, as introduced in [1]. We summarize their functions here for reference.
  • $\sigma$ (Stripping): Removes learned biases inducing collapse.
  • $\alpha$ (Label-Free Abstraction): Projects to geometric relations without labels.
  • $\rho$ (Positioning): Assigns contextual coordinates (implements CIT).
  • $\iota$ (Invariance): Identifies structure stable across contexts.
  • $\delta$ (Dampening): Reduces activation differentials.
  • $\tau$ (Deferred Resolution): Postpones projection to output boundary.
  • $\kappa$ (CPP Integration): Preserves contradictory states with conflict tags.
  • $\pi$ (Persistence): Maintains state across turns with decay.
The composition $\pi\circ\kappa\circ\tau\circ\delta\circ\iota\circ\rho\circ\alpha\circ\sigma$ defines the full pipeline, with $\phi$ (this paper) providing the input transformation from text to initial state.

### v3

@@ Abstract
On a test set of 68 ambiguous sentences, the resulting states preserve interpretive multiplicity: using hybrid extraction we obtain mean state entropy $H=1.087$ bits across ambiguity categories, compared to $H=0$ for collapse-based baselines that commit to a single interpretation. We additionally instantiate the rule-based conflict detector for Japanese markers (kedo, kamoshirenai, etc.) to illustrate cross-lingual portability of the conflict detection stage. This framework extends Non-Resolution Reasoning (NRR) by providing the missing algorithmic bridge between text and the NRR state space, enabling architectural collapse deferment in LLM inference. Design principles for state-to-state transformations are detailed in Appendix D, with empirical validation on 580 test cases (180 single states, 200 contradictory pairs, 200 temporal pairs) demonstrating 0% collapse for principle-satisfying operators versus up to 17.8% for violating operators.
Implementation: A reference implementation accompanying this work is available at https://github.com/kei-saito-research/nrr-phi.
@@ 1 Introduction
This paper is one module in an ongoing Non-Resolution Reasoning (NRR) research program, not a closed endpoint. Companion papers cover foundational theory, implementation strategy, and cross-domain validation with distinct scopes. A series-level Program Map and update status will be maintained in repository documentation.
@@ 1.2 Non-Resolution Reasoning
Non-Resolution Reasoning (NRR) [1] offers an alternative paradigm that treats ambiguity preservation as a valid computational mode. NRR introduced the state space $\mathcal{S}=\{(v_{i},c_{i},w_{i},m_{i})\}$ where multiple interpretations coexist with independent activation weights, and demonstrated that “NRR-lite” implementations (k=2k{=}2 separate embeddings with context gate) maintain high entropy ($H=0.91$ bits, near-maximum $1.0$) at ambiguous turns while baseline models collapse ($H=0.15$ bits).
@@ 1.4 Paper Organization
Section 2 reviews related work on uncertainty in LLM inference. Section 3 presents the formal framework for text-to-state mapping. Section 4 provides algorithms and implementation details. Section 5 reports experimental results. Section 6 briefly illustrates application to psychological support contexts. Section 7 discusses limitations and future work. Section 8 concludes. Appendix D details the design principles and definitions of the NRR transition operators and metadata functions that transform states while preserving non-collapse.
@@ 2.2 Multi-Hypothesis Approaches
Beam search and diverse decoding [7] maintain multiple candidates during generation but ultimately select a single output. Mixture-of-Experts architectures [8, 9] route inputs to specialized subnetworks but still produce collapsed outputs.
@@ 2.4 Relation to NRR Framework
This paper builds directly on NRR-Core [1]. That work established the state space formalism and demonstrated entropy preservation through functional verification (NRR-lite achieving $H=0.91$ bits vs. baseline $H=0.15$ bits). NRR defines a transition-oriented core operator set ($\delta,\tau,\kappa,\pi$), with item-level strengthening used in implementations (e.g., NRR-IME), and additional metadata functions ($\alpha,\rho,\iota$) for representation management (see Appendix D for summary).
@@ 4.2 NRR Processing Pipeline
In this pipeline, $\tau$ is intentionally a no-op on weights: it marks an explicit defer/hold decision instead of forcing an update.
@@ 5.4 Results: Rule-Based Extraction
Note that Theorem 1 requires $|\mathcal{I}(T)|>1$ extractable interpretations. For epistemic, lexical, and structural categories, rule-based methods cannot extract multiple interpretations ($|\mathcal{I}|=1$), so the theorem’s precondition is not met. The $H=0$ results above are not counterexamples but rather motivate the hybrid $\phi$ that combines rule-based and LLM-based extraction.
@@ 5.5 Results: LLM-Based Extraction
To address the limitations of rule-based extraction, we evaluated LLM-based interpretation extraction using three models: ChatGPT, Gemini, and Claude. [^1] Experiments used free-tier web interfaces (January 2026); exact model build IDs were not exposed. Each model was prompted to enumerate all possible interpretations for 28 English sentences (8 epistemic + 10 lexical + 10 structural). The full prompts and raw model outputs are available in the project repository (see prompts/).
@@ 5.6 Combined Results
Overall 68 2.24 1.087 Hybrid
Table 4: Combined results using hybrid extraction. All categories achieve $H>0$, with overall mean entropy $H=1.087$ bits. The overall row is computed as the weighted mean across all 68 sentences, where LLM-based $H$ values represent the three-model average for each sentence.
@@ Appendix C Sample Outputs
Appendix D NRR Operator Design Principles and Definitions
This appendix details design principles and formal definitions for NRR’s state-transition machinery. We distinguish (i) core transition operators used for state evolution and (ii) metadata functions used for representation management. While the $\phi$ mapping (main text) handles text-to-state transformation, these components govern state-to-state processing.
D.1 Design Principles for Non-Collapsing Operators
Transforming NRR states risks destroying the multiplicity that NRR preserves. We identify four principles that operators must satisfy to maintain non-collapse. These principles derive from the structure of the NRR state space itself.
D.1.1 Principle 1: Relative Structure Preservation
Principle 1 (Relative Structure Preservation).
An operator must preserve the proportional relationships between interpretation weights. If $w_{i}/w_{j}=r$ before the operation, then $w^{\prime}_{i}/w^{\prime}_{j}\approx r$ after.
Rationale: Entropy depends on relative, not absolute, weight values. Transformations that preserve weight ratios maintain the entropy structure. Note that ratio preservation is a sufficient condition for non-collapse, but not strictly necessary—operators such as dampening ($\delta$) preserve non-collapse by increasing entropy, even when ratios are not strictly preserved.
Implementation: Use multiplicative transformations (e.g., $w^{\prime}_{i}=\lambda w_{i}$) or proportional adjustments (e.g., $w^{\prime}_{i}=w_{i}-b\cdot w_{i}/\max w$) rather than uniform additive changes.
D.1.2 Principle 2: Scale Equivariance
Principle 2 (Scale Equivariance).
An operator’s effect should be independent of the absolute scale of weights. Scaling all weights by a constant should yield the same entropy change.
Rationale: Since weights are non-normalized, absolute magnitudes are arbitrary. Operations should depend only on relative structure.
Implementation: Design operators such that $\mathcal{O}(\lambda\mathcal{S})\sim\lambda\mathcal{O}(\mathcal{S})$ for scalar $\lambda>0$. This property is verifiable analytically from the operator definition.
D.1.3 Principle 3: Contradiction Non-Destruction
Principle 3 (Contradiction Non-Destruction).
When merging states with conflicting interpretations, preserve both rather than eliminating one. Contradictions are information, not errors.
Rationale: This implements the Contradiction-Preservation Principle from NRR foundations. Contradictions are structure, not error; destroying them collapses information.
Implementation: Union with conflict tagging rather than replacement or selection.
D.1.4 Principle 4: Temporal Persistence
Principle 4 (Temporal Persistence).
Interpretations from previous states should persist (with decay) rather than being overwritten, enabling reactivation when context shifts.
Rationale: Context may shift to favor previously dormant interpretations. Complete overwriting destroys information that may become relevant.
Implementation: Decayed union of current and previous states.
D.1.5 The Information Preservation Law
The principles converge on a single quantitative criterion.
Definition 10 (Information Loss).
The information loss induced by operator $\mathcal{O}$ is:
$L(\mathcal{O},\mathcal{S})=H(\mathcal{S})-H(\mathcal{O}(\mathcal{S}))$ (15)
Positive $L$ indicates collapse; negative $L$ indicates information gain.
Definition 11 (Information Preservation Law).
An operator $\mathcal{O}$ satisfies the $\epsilon$-preservation law if:
$H(\mathcal{O}(\mathcal{S}))\geq H(\mathcal{S})-\epsilon$ (16)
for all states $\mathcal{S}$ and small threshold $\epsilon$. Equivalently, $L(\mathcal{O},\mathcal{S})\leq\epsilon$.
Operators satisfying Principles 1–4 will satisfy the preservation law; operators violating them risk violation.
D.2 Operator Definitions
We define four core transition operators that have been empirically validated to preserve non-collapse (§D.6), plus one state-level calibration check and three metadata functions.
D.2.1 Core Transition Operators (Empirically Validated)
The following core transition operators have been validated on 180 single states, 200 contradictory pairs, and 200 temporal pairs across three LLMs.
Dampening ($\delta$).
Compresses weights toward the mean, reducing dominance of any single interpretation:
$\delta(\mathcal{S})=\{(v_{i},c_{i},w_{i}(1-\lambda)+\bar{w}\lambda)\}$ (17)
where $\bar{w}=\frac{1}{n}\sum_{j}w_{j}$. Dampening does not strictly preserve weight ratios; rather, it satisfies the spirit of Principle 1 by preventing undue suppression of minority interpretations and actually increases entropy, thereby satisfying the $\epsilon$-preservation law. Empirically achieves 0% collapse across all tested parameter settings.
State-level Calibration ($\sigma_{\mathrm{state}}$).
Applies a uniform multiplicative scaling to all weights (state-level calibration check):
$\sigma_{\mathrm{state}}(\mathcal{S})=\left\{\left(v_{i},c_{i},w_{i}-b\cdot\frac{w_{i}}{\max_{j}w_{j}}\right)\right\}$ (18)
This operation multiplies all weights by the same factor $\left(1-b/\max_{j}w_{j}\right)$. Therefore, after normalization for entropy evaluation, the induced probability distribution is unchanged and $\Delta H=0$. We retain this as a scale-equivariance calibration operator, not as an entropy-changing update rule. This $\sigma_{\mathrm{state}}$ is separate from item-level strengthen operators used in NRR-IME/Universal.
Deferred Resolution ($\tau$).
Identity during processing; resolution only at output boundary:
$\tau(\mathcal{S})=\mathcal{S}$ (19)
Satisfies all principles by doing nothing until output is required. Serves as baseline in empirical validation.
CPP Integration ($\kappa$).
Implements Principle 3 (Contradiction Non-Destruction):
$\kappa(\mathcal{S},\mathcal{S}^{\prime})=\mathcal{S}\cup\mathcal{S}^{\prime}\text{ with conflict tags}$ (20)
Contradictory information is preserved, not eliminated. Validated on 200 contradictory pairs with 0% collapse.
Persistence ($\pi$).
Implements Principle 4 (Temporal Persistence):
$\pi(\mathcal{S}_{t},\mathcal{S}_{t-1})=\mathcal{S}_{t}\cup\{(v_{i},c_{i},\gamma w_{i}):(v_{i},c_{i},w_{i})\in\mathcal{S}_{t-1}\}$ (21)
Previous interpretations persist with decayed activation ($\gamma\in(0,1)$ is the decay coefficient). Validated on 200 temporal pairs with 0% collapse.
D.2.2 Metadata Functions (Representation Utilities)
The following three functions preserve weights by construction and primarily augment representation metadata rather than perform transition dynamics. They are included as extensible utilities.
Positioning ($\rho$).
Assigns temporal coordinates to distinguish identical tokens across contexts:
$\rho(\mathcal{S})=\{(v_{i},(c_{i},t),w_{i})\}$ (22)
Implements contextual identity tracking without affecting weights. Preserves entropy by construction since no weights are modified.
Abstraction ($\alpha$).
Augments state with relational structure:
$\alpha(\mathcal{S})=\{(v_{i},c_{i},w_{i},R_{i})\}$ (23)
where $R_{i}=\{d(v_{i},v_{j}):j\neq i\}$ captures geometric relations. Preserves entropy by construction since no weights are modified.
Invariance ($\iota$).
Marks interpretations originating from the same symbol across multiple contexts:
$\iota(\mathcal{S}_{1},\ldots,\mathcal{S}_{n})=\bigcup_{j=1}^{n}\mathcal{S}_{j}\text{ with cross-context ID}$ (24)
where each interpretation $s_{i}\in\mathcal{S}_{j}$ receives metadata marking it as derived from the same underlying symbol (e.g., the word “bank” appearing in financial vs. riverbank contexts). By construction, $\iota$ retains all interpretations from all input contexts: $|\iota(\mathcal{S}_{1},\ldots,\mathcal{S}_{n})|=\sum_{j}|\mathcal{S}_{j}|$, preserving or increasing entropy.
D.2.3 Non-Destructive Projection
For output generation:
$\Pi(\mathcal{S})=\arg\max_{i}w_{i},\quad\mathcal{S}^{\prime}=\mathcal{S}$ (25)
The system produces singular output while preserving internal state. Note that $\Pi$ is an output mechanism, not a processing operator; it is not subject to the design principles, since its purpose is precisely to select a single interpretation for external communication.
D.3 Theoretical Properties
D.3.1 Operator–Data Structure Alignment
Different components require different data structures for validation. Weight-transforming transitions ($\delta$, $\sigma_{\mathrm{state}}$, $\tau$) operate on individual states and are naturally validated on single states. In contrast, $\kappa$ (contradiction integration) and $\pi$ (temporal persistence) are binary operators that take two states as input—validating them on single states would be meaningless, as it would fail to exercise their core function. Metadata functions that neither modify weights nor filter interpretations ($\alpha$, $\rho$, $\iota$) preserve entropy by construction.
D.3.2 Composition Behavior
Since each principle-satisfying operator individually satisfies $H(\mathcal{O}(\mathcal{S}))\geq H(\mathcal{S})-\epsilon$, a chain of $k$ such operators accumulates at most $k\epsilon$ information loss. For the transition components in this appendix, this bound is supported analytically: $\sigma_{\mathrm{state}}$ applies an identical multiplicative factor to all weights, so $\Delta H=0$ exactly (the normalized distribution is invariant); $\delta$ moves weights toward the mean, which by the strict concavity of Shannon entropy guarantees $\Delta H\geq 0$; $\tau$ is the identity, so $\Delta H=0$ trivially.
For $\epsilon=0.1$ bits, a three-component chain $\delta\circ\tau\circ\rho$ is bounded by $\Delta H\leq 0.3$ bits in the worst case. In practice, since $\delta$ increases entropy and $\rho$ leaves it unchanged, the actual cumulative loss is substantially smaller.
D.3.3 Analytical Verification of Scale Equivariance
Principle 2 (Scale Equivariance) requires that scaling all weights by a constant $c>0$ does not change the entropy outcome. Since $H(\mathcal{S})$ depends only on normalized probabilities $p_{i}=w_{i}/\sum_{j}w_{j}$, and scaling all $w_{i}$ by $c$ leaves $p_{i}$ invariant, we verify that each operator commutes with scaling:
  • $\sigma_{\mathrm{state}}$ (State-level calibration): $\sigma_{\mathrm{state}}(c\mathcal{S})_{i}=cw_{i}-b\cdot cw_{i}/c\max w=c(w_{i}-b\cdot w_{i}/\max w)=c\cdot\sigma_{\mathrm{state}}(\mathcal{S})_{i}$. Exact scale equivariance.
  • $\delta$ (Dampening): $\delta(c\mathcal{S})_{i}=cw_{i}(1-\mu)+c\bar{w}\mu=c[w_{i}(1-\mu)+\bar{w}\mu]=c\cdot\delta(\mathcal{S})_{i}$. Exact scale equivariance.
  • $\tau$ (Identity): Trivially equivariant.
  • $\kappa$ (CPP): Union of sets; scaling both input states by $c$ scales all weights in the output by $c$. Equivariant.
  • $\pi$ (Persistence): $\pi(c\mathcal{S}_{t},c\mathcal{S}_{t-1})$ scales all weights (including decayed ones) by $c$. Equivariant.
  • $\rho$, $\alpha$, $\iota$: These operators do not modify weights, so scale equivariance is trivially satisfied.
All core transitions and metadata functions satisfy Principle 2 analytically.
D.4 Operator–Principle Correspondence
Table 6 summarizes which principles each operator satisfies. P1 = Relative Structure Preservation, P2 = Scale Equivariance, P3 = Contradiction Non-Destruction, P4 = Temporal Persistence. Operators marked with have been empirically validated (§D.6).
Operator P1 P2 P3 P4
Empirically validated core transitions:
$\delta$ (Dampening) $\approx*$
$\sigma_{\mathrm{state}}$ (Calibration)
$\tau$ (Hold/Defer)
$\kappa$ (CPP)
$\pi$ (Persistence)
Metadata functions (representation utilities):
$\rho$ (Positioning)§
$\alpha$ (Abstraction)
$\iota$ (Invariance)
Table 6: Operator–principle correspondence. Empirically validated operators (§D.6). $\rho$ only adds context metadata and does not modify weights, so entropy preservation holds by construction; empirically confirmed across all validation runs (§D.6). $\delta$ does not strictly preserve weight ratios (P1) but satisfies the $\epsilon$-preservation law by increasing entropy. $\iota$ satisfies P1 by retaining all interpretations from all contexts.
D.5 Pipeline Composition
For implementation, a compact transition chain $\pi\circ\kappa\circ\delta\circ\tau$ is typically sufficient, with $\phi$ (main text) providing the input transformation from text to initial state. Optional metadata functions ($\rho,\alpha,\iota$) can be inserted where context tracking or cross-context labeling is required. Algorithm 2 shows this compact pattern with optional positioning. Each validated transition preserves the interpretive multiplicity established by $\phi$, enabling non-collapsing inference through repeated updates.
D.6 Empirical Validation
We validate that principle-satisfying operators achieve lower collapse rates than principle-violating alternatives. Each operator is validated on the data structure appropriate to its function.
D.6.1 Methodology
Data: Three datasets were constructed using the NRR-Phi $\phi$ mapping procedure on a newly constructed sentence set, entirely distinct from the 68-sentence test set in the main text:
  1. 1.
    Single states ($n=180$): 60 ambiguous sentences (40 epistemic, 20 lexical; English and Japanese) processed by three LLMs (ChatGPT, Gemini, Claude), yielding 180 states. Each state contains 2–4 interpretations with confidence weights.
  2. 2.
    Contradictory pairs ($n=200$): 150 pairs from the same sentence processed by different models (capturing inter-model disagreement; 50 sentences $\times$ $\binom{3}{2}=3$ model pairs, selected from the 60 sentences by excluding 10 with near-identical model outputs), plus 50 pairs from ambiguous words in different contexts (capturing intra-word polysemy).
  3. 3.
    Temporal pairs ($n=200$): 150 two-turn dialogue pairs (where context evolves across turns), plus 50 context-evolution pairs (where the same sentence is reinterpreted under shifted context).
Metric: Violation rate—percentage of states where $H(\mathcal{S}^{\prime})<H(\mathcal{S})-\epsilon$ with $\epsilon=0.1$ bits. This threshold corresponds to 10% information loss relative to binary states ($H_{\max}=1.0$ bit).
Operator–data matching: Weight-transforming transitions ($\delta$, $\sigma_{\mathrm{state}}$, $\tau$) are tested on single states. The integration operator ($\kappa$) is tested on contradictory pairs. The persistence operator ($\pi$) is tested on temporal pairs. This matching ensures each operator is validated on the data structure it is designed to process.
Comparison: For weight-transforming operators, we compare:
  • Principle-violating: Uniform subtraction ($\delta$ v1: $w^{\prime}_{i}=w_{i}-b$)
  • Principle-satisfying: Proportional transformation ($\delta$ v2: $w^{\prime}_{i}=w_{i}(1-\lambda)+\bar{w}\lambda$; $\sigma_{\mathrm{state}}$: $w^{\prime}_{i}=w_{i}-b\cdot w_{i}/\max w$)
Parameter sweep: $\delta$ v1 was tested at three subtraction magnitudes ($b\in\{0.05,0.10,0.20\}$), $\delta$ v2 at five compression rates ($\lambda\in\{0.1,0.2,0.3,0.4,0.5\}$), and $\sigma_{\mathrm{state}}$ at four bias levels ($b\in\{0.05,0.10,0.15,0.20\}$). This yields 2,740 total operator–state measurements.
D.6.2 Results
Figure 4: Entropy change by operator type (representative parameter $b=0.10$ shown for $\delta$ v1). The principle-violating uniform operator ($\delta$ v1, shown in red) collapses information in 6.1% of cases at $b=0.10$, falling below the $\Delta H=-\epsilon=-0.1$ threshold (red dashed line); other settings yield 1.7% ($b=0.05$) and 17.8% ($b=0.20$). All principle-satisfying operators ($\delta$ v2, $\sigma_{\mathrm{state}}$, $\tau$, $\kappa$, $\pi$, shown in green) achieve 0% collapse. Operators $\kappa$ and $\pi$ substantially increase entropy through integration and persistence mechanisms. The near-zero bars for $\sigma_{\mathrm{state}}$ and $\tau$ are expected: $\sigma_{\mathrm{state}}$ is a calibration check, and $\tau$ is an intentional hold step.
Figure 5: Analysis across ambiguity categories (epistemic, lexical). (A) Mean entropy by category with standard deviation error bars; both categories maintain high multiplicity ($H>1.7$ bits), approaching the theoretical maximum for four interpretations ($H_{\max}=2.0$ bits). (B) Average number of interpretations extracted per state; dashed line indicates minimum target ($|\mathcal{S}|\geq 2$). (C) Entropy increase from dampening operator ($\delta$ v2) relative to pre-dampening state, demonstrating consistent preservation across both categories.
Figure 4 shows entropy change by operator type. The principle-violating uniform operator ($\delta$ v1) collapses information at a rate that increases with subtraction magnitude: 1.7% at $b=0.05$, 6.1% at $b=0.10$, and 17.8% at $b=0.20$ (32 of 180 states). In contrast, all principle-satisfying operators achieve 0% collapse across all parameter settings.
Table 7 presents the complete results.
Operator Data $n$ Violations Mean $\Delta H$
$\delta$ v1 ($b=0.05$) single 180 3 (1.7%) 0.012-0.012
$\delta$ v1 ($b=0.10$) single 180 11 (6.1%) 0.030-0.030
$\delta$ v1 ($b=0.20$) single 180 32 (17.8%) 0.086-0.086
$\delta$ v2 ($\lambda=0.1$) single 180 0 (0%) +0.012+0.012
$\delta$ v2 ($\lambda=0.3$) single 180 0 (0%) +0.031+0.031
$\delta$ v2 ($\lambda=0.5$) single 180 0 (0%) +0.044+0.044
$\sigma_{\mathrm{state}}$ ($b=0.10$) single 180 0 (0%) $0.000$
$\sigma_{\mathrm{state}}$ ($b=0.20$) single 180 0 (0%) $0.000$
$\tau$ (identity, baseline) single 180 0 (0%) $0.000$
$\kappa$ (CPP) contra. pairs 200 0 (0%) +0.881+0.881
$\pi$ (persistence) temporal pairs 200 0 (0%) +0.922+0.922
Table 7: Experimental validation results (representative parameter values shown; all omitted settings also achieved 0% violations). Principle-violating operators ($\delta$ v1) show increasing collapse rates with subtraction magnitude. All principle-satisfying operators achieve 0% violations across all parameter settings tested. $\kappa$ and $\pi$ show large positive $\Delta H$ because they merge interpretations from two states, increasing multiplicity. $\sigma_{\mathrm{state}}$ applies an identical multiplicative factor $({1-b/\max w})$ to all weights, leaving the normalized distribution invariant; $\Delta H=0$ is analytically exact (verified computationally). Similarly, $\tau$ (identity) yields $\Delta H=0$ by mathematical necessity.
Figure 5 presents analysis across the two ambiguity categories from the single-state dataset. Epistemic ambiguities maintain mean entropy $H=1.81$ bits (SD =0.49=0.49) with an average of 3.76 interpretations per state. Lexical ambiguities show $H=1.75$ bits (SD =0.33=0.33) with 3.68 interpretations. Both categories demonstrate that the $\phi$ mapping reliably extracts multiple distinct interpretations.
D.6.3 Key Findings
  1. 1.
    Collapse rate scales with violation magnitude: $\delta$ v1 violation rate increases monotonically from 1.7% ($b=0.05$) to 17.8% ($b=0.20$), confirming that larger uniform subtractions disproportionately suppress weaker interpretations.
  2. 2.
    Dampening increases entropy: $\delta$ v2 shifts distributions toward uniformity, raising $H$ by an average of 0.03 bits at $\lambda=0.3$. This increase is analytically guaranteed by the strict concavity of Shannon entropy (see Appendix D) and confirmed empirically.
  3. 3.
    State-level calibration is entropy-invariant: $\sigma_{\mathrm{state}}$ yields mean $\Delta H\approx 0$ bits by design, providing a scale-equivariance check without changing the normalized interpretation distribution.
  4. 4.
    Integration and persistence add information: $\kappa$ and $\pi$ increase $H$ by +0.88+0.88 and +0.92+0.92 bits respectively, by merging interpretations from paired states.
  5. 5.
    Zero violations for all principle-satisfying operators: Across 180 single states and 400 pairs, no principle-satisfying operator exhibited $\Delta H<-0.1$ bits under any parameter setting tested.